Circular No. 7326

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gg IN MCG +7-8-11 G. M. Hurst, Basingstoke, England, reports the discovery by Tom

Boles, Wellingborough, of a supernova (mag ∼ 16.8) on an unfiltered CCD frame taken on Nov. 6.965 UT in the course of the U.K. Nova/Supernova Patrol with a 0.36-m reflector; the new object is lcoated at α = 3^{h}27^{m}58^{s}.9, δ = +39^{o}54′18′′.4 (equinox 2000.0). M. Schwartz, Cottage Grove, OR, obtained a CCD image showing SN 1999gg on Nov. 23, but the new object is not visible to mag ∼ 18 on an image taken by Schwartz on 1998 Nov. 1.

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that spectra of SN 1999gg, obtained by E. Falco on Dec. 2.28 and 3.25 UT with the F. L. Whipple Observatory (FLWO) 1.5-m telescope (+ FAST spectrograph), show it to be a type-II supernova. The spectrum is heavily contaminated by the host galaxy, but a broad P-Cyg Hα profile is visible, as well as broad Hβ absorption. The photospheric expansion velocity of the supernova, measured from the Hα absorption minimum, is 5100 km/s. Cross-correlation of the host galaxy spectrum with an emission-line template yields a recession velocity of 4350 ± 20 km/s. An image of SN 1999gg, taken by A. Bragg on Nov. 30.29 with the FLWO 1.2-m telescope (+ 4Shooter), yields precise position end figures for the supernova: 58^{s}.84, 19′′.0 (equinox 2000.0, uncertainty ± 0′′.2), which is 4′′.7 west and 3′′.3 north of the galaxy nucleus.

SUPERNOVA 1999gb IN NGC 2532 Jha, Challis, and Kirshner also report: “A spectrum of SN 1999gb (cf.

IAUC 7316), obtained by Falco on Nov. 30.45 UT with the FLWO 1.5-m telescope (+ FAST spectrograph), reveals that it is likely a type-IIn supernova at an early epoch. The spectrum exhibits a very blue continuum with Hα emission, the line profile consisting of very narrow emission (FWHM < 300 km/s) on a broader base (FWHM = 2800 km/s). The redshift of the host galaxy, measured via cross-correlation with an emission-line template is 5270 ± 15 km/s.”

V1494 AQUILAE Visual magnitude estimates: Dec. 2.95 UT, 3.7 (J. Rao, Yonkers, NY);

3.043, 4.1 (S. Giovanardi, New York City, NY); 3.38, 4.1 (H.-l. Tao, Raohe, Heilongjiang, China); 3.46, 3.6 (W.-f. Song, Jingzhou, China); 3.670, 4.1 (K. Hornoch, Lelekovice, Czech Republic); 3.68, 4.0 (L. Kiss, Szeged, Hungary); 3.702, 4.3 (Hornoch).

1999 December 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green